OpenStudy (anonymous):

solve this system using elimination 3x-2y=0 4x+2y=14

OpenStudy (anonymous):

3x-2y=0 (*2) -> 6x-4y=0 4x+2y=14 (*2) -> 8x+4y=28 14x=28 x=2 3(2)-2y=0 6-2y=0 +6 +6 -2y=6 y= -3